OpenStudy (surryyy):

The area of the triangle shown below is 60 cm2. Find the value of x. answers: 6 12 18 36

OpenStudy (surryyy):



OpenStudy (midhun.madhu1987):

Area of a triangle = 1/2 * base * height

Directrix (directrix):

A = (1/2) * 10 * (2x) Substitute 60 for area. So, 60 = (1/2) * 10 * (2x) Solve for x @surryyy

OpenStudy (midhun.madhu1987):

 Area = 60 so find x